OpenStudy (anonymous):

You deposit $100 into a savings plan at the end of each year. The interest rate is 6% compounded annually. Find the value of the annuity after 7 years. Do not round until the final answer. Then, round to the nearest cent

OpenStudy (anonymous):

compounding interest formula A=P(1+r/t)^rt, in this case r=.07 and t=# of years; a series of these can be used to find the answer so... A=100(1.07)^7+100(1.07)^6+...+100(1.07)^2+100(1.07)^1=765.4021

OpenStudy (anonymous):

A $2,506.05 B $697.53 C $839.38 D $236.42

OpenStudy (anonymous):

Those are the options @coureges

OpenStudy (anonymous):

my interest rate was off, 839.38

OpenStudy (wolf1728):

This would be called an annuity and the formula for calculating an annuity is attached. What bothered me about this problem was the lousy way in which it was worded. It seems it is asking for a full seven years of interest. No. It seems that the time begins with the first deposit of $100.00 and so interest is collected only for 6 total years and the interest earned on the final $100.00 is zero. Years Total 6 141.85 5 133.82 4 126.25 3 119.10 2 112.36 1 106.00 0 100.00 839.38